OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half. y - 6 = one half(x + 3) y = one halfx - 6 y + 3 = one half(x - 6) x - 2y = 12

OpenStudy (anonymous):

@cj49 @heavenfun13

OpenStudy (cj49):

m=1/2 y=1/2x+b x,y=6,-3 -3=1/2(6)+b -3=3+b b=-6 y=1/2x-6